If a sheet containing a single slit is heated (without damaging it) and therefore expands, what happens to the angular location of the first-order diffraction minimum?

It moves toward the centerline.
It moves away from the centerline.
It doesn't change.
Answer #1

First order diffraction is given by:

(m + 1/2)*lambda = d*sin theta

m = a

3*lambda/2 = d*sin theta

theta = arcsin (3*lambda/(2*d))

Now if we increase temperature, then we know that length of sheet will increase, So if d will increase, then from above equation theta will decrease.

So It move toward the centerline.

Please Upvote.